# arXiv:2603.00301v1[math.NT] 27 Feb 2026

### 27 Feb 2026 v1

@@ Abstract.
For a real quadratic field $\mathbb{Q}(\sqrt{d})$, we study the norm-form energy $N=S_{\zeta}^{2}-d\cdot S_{L}^{2}$, where $S_{\zeta}$ and $S_{L}$ are Lorentzian-weighted zero sums with $w(\rho)=2/(\tfrac{1}{4}+\gamma^{2})$. We prove three main results. (1) Spacelike spectral data: $N<0$ unconditionally for all squarefree $d>1$, as a consequence of a low-lying zero dominance theorem proved via explicit zero-counting. (2) Effective density bound: at each verified truncation level $M$, $\operatorname{dens}\{N>0\}\leq 2\|f_{S_{L}^{(M)}}\|_{\infty}\cdot(W_{1}(\zeta)/\sqrt{d}+\epsilon_{M})$, established unconditionally via Jacobi–Anger resonance analysis. (3) Exact asymptotic: under the computationally verified hypothesis that the infinite resonance lattice $\Lambda_{\infty}$ has finite rank (verified for $M\leq 20$, where rank =0=0), the sharp asymptotic $\operatorname{dens}\{N>0\}=C(d)/\sqrt{d}+o(1/\sqrt{d})$ holds. For $d=5$, $C(5)=2\,f_{S_{L}}(0)\cdot\mathbb{E}[|S_{\zeta}|]=0.1191$; the constant depends on $d$ through the zeros of $L(s,\chi_{d})$, and $C(d)=O(1/\log d)$ as $d\to\infty$.
@@ 1.2. Logical structure and the GSH barrier
Thm 4.16: Density vanishing Hypothesis 4.4(ii) $O(1/\sqrt{d})$ Jessen–Wintner
Thm 4.18: Euler product density nothing $O(1/\sqrt{d})$ $\mathbb{Q}$-indep. of $\log p$
@@ 2.1. The Lorentzian weight
Since every nontrivial zero of $\zeta(s)$ has $|\gamma|\geq 14.13>1/(2\sqrt{3})$, all estimates using the on-line weight $w(\rho)=2/(\tfrac{1}{4}+\gamma^{2})$ are unconditional upper bounds. For $L$-function zeros, values $\gamma^{\prime}<1/(2\sqrt{3})\approx 0.289$ are theoretically possible (and expected for large conductors), but irrelevant here: the general weight formula (5) applies unconditionally to all zeros regardless of $\gamma^{\prime}$, and the on-line formula is used only as an upper bound via Proposition 2.2.
@@ 3.1. Low-lying zero dominance
We establish the inequality for all fundamental discriminants by combining Theorem 3.2 with verified computations for two small-conductor exceptions. Throughout, zeros are counted in the upper half-plane only; the weight $w(\rho)=2/(\tfrac{1}{4}+\gamma^{2})$ represents conjugate pairs.
@@
where $\ell=\log(q\cdot 13/(2\pi))$. For $q=12$, this evaluates to $N(11,\chi_{d})\geq 3.80$; since $N(11,\chi_{d})$ is an integer, this implies at least four zeros below height $11$, and the bound increases with $q$. Each of these at least four zeros has imaginary part at most $11$, so each contributes weight at least $w(11)=2/(\tfrac{1}{4}+121)=2/121.25=0.01649$. Therefore
$S_{L}\geq 4\cdot w(11)=8/121.25=0.06598>0.04871=S_{\zeta}^{*}.$
Step 2: $d=5$ (conductor $q=5$). The BMOR bound gives only $N(11,\chi_{5})\geq 1.40$, which does not guarantee three zeros. We therefore verify the inequality directly. Computing the Hardy $Z$-function $Z_{\chi_{5}}(t)=e^{i\theta(t)}L(\tfrac{1}{2}+it,\chi_{5})$ and scanning for sign changes on $(0,12)$ yields two zeros below height $11$: $\gamma_{1}^{\prime}=6.6485$ and $\gamma_{2}^{\prime}=9.8314$, both verified to satisfy $|L(\tfrac{1}{2}+i\gamma^{\prime},\chi_{5})|<10^{-15}$. (Any additional zeros below height $11$, if they exist, would only increase $S_{L}$.) Therefore
$S_{L}\geq w(\gamma_{1}^{\prime})+w(\gamma_{2}^{\prime})=\frac{2}{0.25+44.20}+\frac{2}{0.25+96.66}=0.04499+0.02064=0.06563>S_{\zeta}^{*}.$
@@
$\gamma_{1}^{\prime}=4.8999739970070365010,\quad\gamma_{2}^{\prime}=7.6284288417693978341,\quad\gamma_{3}^{\prime}=10.806588163861712014,$
@@
$S_{L}\geq w(\gamma_{1}^{\prime})+w(\gamma_{2}^{\prime})+w(\gamma_{3}^{\prime})=\frac{2}{0.25+24.010}+\frac{2}{0.25+58.193}+\frac{2}{0.25+116.782}=0.08246+0.03421+0.01710=0.13377>S_{\zeta}^{*}.$
@@
The margin $(S_{L}-S_{\zeta}^{*})/S_{\zeta}^{*}$ is smallest in the generic case: for $q=12$, the BMOR bound guarantees at least four zeros below height $11$, giving $S_{L}\geq 4\cdot 2/(0.25+121)=0.06598$, a margin of $(0.06598-0.04871)/0.04871=35.5\%$ over $S_{\zeta}^{*}$. The margin is largest for $d=2$, where three directly computed zeros give $S_{L}\geq 0.13375$, a margin of $174.6\%$. The two exceptional cases $d\in\{2,5\}$ are verified by direct computation of $L$-function zeros via the Hardy $Z$-function method (see Appendix C for the $d=2$ data).
@@
For most discriminants, the first zero alone gives $w(\gamma_{1}^{\prime})>S_{\zeta}^{*}$. The exception is $d=5$ ($\gamma_{1}^{\prime}=6.6485$), where $w(\gamma_{1}^{\prime})=0.04499<S_{\zeta}^{*}=0.04871$ and the second zero is required.
@@ 3.3. The negativity theorem for general characters
The proof follows the same structure as Theorem 3.1. For $q\geq 4$, Theorem 3.2 guarantees a zero below height $T=10.5$, giving $q\cdot w(\gamma_{1}^{\prime})\geq 4\cdot 2/110.50=0.0724>S_{\zeta}^{*}$. For $q=3$, the BMOR bound at $T=15.7$ guarantees at least three zeros, each below height $15.7$, so $q\cdot S_{L(\chi)}\geq q\cdot 3\cdot w(15.7)=3\cdot 3\cdot 2/(0.25+15.7^{2})=18/246.74=0.0730>S_{\zeta}^{*}$. ∎
@@
By (6), $S_{\zeta}\leq S_{\zeta}^{*}$. For $q\geq 4$, the BMOR bound guarantees a zero of $L(s,\chi)$ below height $T(q)\leq 10.5<11$, so $S_{L(\chi)}\geq w(11)=2/121.25$. For $q=3$, the BMOR bound at $T=15.7$ gives $S_{L(\chi)}\geq 3\cdot 2/(0.25+15.7^{2})=6/246.74>2/121.25$. In both cases, $q\cdot S_{L(\chi)}\geq 2q/121.25$, and $E_{\chi}(0)=S_{\zeta}-q\cdot S_{L(\chi)}\leq S_{\zeta}^{*}-2q/121.25$. Since $2\cdot 3/121.25=0.04948>S_{\zeta}^{*}=0.04871$, this is negative for all $q\geq 3$. ∎
@@ 3.5. Robustness
The negativity and spacelike results are robust to failure of RH. If the Riemann Hypothesis fails for $\zeta(s)$, off-line zeros have smaller Lorentzian weight by Proposition 2.2, decreasing $S_{\zeta}$. If RH fails for $L(s,\chi_{d})$ via a Siegel zero $\rho_{0}=\beta_{0}$ with $\beta_{0}$ near $1$, the weight $w(\rho_{0})=2/(\beta_{0}(1-\beta_{0}))$ diverges, increasing $S_{L}$. Both scenarios strengthen the inequalities.
@@ 4.1. Shifted zero sums
Each is a uniformly convergent sum of cosines with incommensurable frequencies (the zero ordinates), which defines a uniformly almost periodic function in the sense of Theorem 4.2. ∎
@@ 4.2. The Grand Simplicity Hypothesis
The unconditional results of this paper (Theorems 3.1, 3.7, 4.9, and 8.5) require neither sub-case.
@@ 4.3. Parseval identity
Assume Hypothesis 4.4 (i) (cross-function disjointness of zero ordinate sets). Then the Cesàro mean of the norm-form energy satisfies
@@
so the disjointness-free identity gives a weaker (less negative) bound on $\langle N\rangle$.
@@
By Theorem 4.2, $\langle S_{\zeta}(\delta)^{2}\rangle=\tfrac{1}{2}W_{2}(\zeta)$ and $\langle S_{L}(\delta)^{2}\rangle=\tfrac{1}{2}W_{2}(L)$, both unconditionally (since all zeros of each function are simple). The cross-term $\langle S_{\zeta}\cdot S_{L}\rangle=\tfrac{1}{2}\sum_{\gamma_{k}=\gamma_{j}^{\prime}}w(\gamma_{k})w(\gamma_{j}^{\prime})$ vanishes when the zero ordinate sets are disjoint. ∎
@@ 4.4. Cesàro spacelike theorem
For $k\geq 2$, set $a_{1}=w(\gamma_{1})=2/(\tfrac{1}{4}+\gamma_{1}^{2})$ (the largest weight in $\{a_{j}\}$) and $b_{1}=w(\gamma_{1}^{\prime})=2/(\tfrac{1}{4}+{\gamma_{1}^{\prime}}^{2})$ (the largest weight in $\{b_{k}\}$). Since $\gamma_{1}^{\prime}<\gamma_{1}$ (established above), we have $b_{1}>a_{1}$; define $r(d)=a_{1}/b_{1}=(\tfrac{1}{4}+{\gamma_{1}^{\prime}}^{2})/(\tfrac{1}{4}+\gamma_{1}^{2})<1$.
@@
$W_{k}(\zeta)=\sum_{j}a_{j}^{k}\leq a_{1}^{k-1}\sum_{j}a_{j}=a_{1}^{k-1}\cdot W_{1}(\zeta).$
Lower bound on $W_{k}(L)$: $W_{k}(L)\geq b_{1}^{k}=(a_{1}/r(d))^{k}$.
Combining: $W_{k}(\zeta)/W_{k}(L)\leq a_{1}^{k-1}W_{1}(\zeta)\cdot r(d)^{k}/a_{1}^{k}=W_{1}(\zeta)\cdot r(d)^{k}/a_{1}$. Since $r(d)<1$, this ratio decays exponentially in $k$. The inequality $W_{k}(\zeta)<d\cdot W_{k}(L)$ holds for all $d\geq 2$ and $k\geq 2$: indeed, $d\cdot W_{k}(L)\geq d\cdot b_{1}^{k}>W_{k}(\zeta)$ because $d\cdot b_{1}>a_{1}$ (as $b_{1}>a_{1}$ and $d>1$) implies $d\cdot b_{1}^{k}>a_{1}^{k-1}b_{1}>a_{1}^{k-1}\cdot\sum_{j}a_{j}\cdot(a_{j}/a_{1})^{k-1}=W_{k}(\zeta)$ for $k\geq 2$, where the last inequality uses $(a_{j}/a_{1})^{k-1}\leq 1$. ∎
@@
for all $T>0$, where $c_{0}=(\pi-1)/(2\pi)$. For $q\geq 17$ and $T=11$, Theorem 3.2 guarantees $N(11,\chi_{d})\geq 5$, giving $W_{2}(L)\geq 5(8/485)^{2}=320/235225$. Therefore $c_{0}W_{2}(L)\geq 320(\pi-1)/(470450\pi)>4.63\times 10^{-4}$, while $(S_{\zeta}^{*})^{2}/d\leq(0.04871)^{2}/14<1.70\times 10^{-4}$, a margin of $2.7\times$ or better for all $d\geq 14$. For the worst-case five-zero configuration (equally spaced in $[6,11]$), the interference constant evaluates to $\mathcal{I}=5.62\times 10^{-3}$, so the full Cesàro variance exceeds the threshold for all $T>2.6$.
@@
For small $T$: as $T\to 0^{+}$, $\langle S_{L}^{2}\rangle_{T}\to W_{1}(L)^{2}\geq(0.06598)^{2}=4.35\times 10^{-3}$, which exceeds $(S_{\zeta}^{*})^{2}/d\leq 1.70\times 10^{-4}$ by a factor of $25.6$. By the same Lipschitz bound used in Case 2 (with $|h^{\prime}(T)|\leq L$), the inequality persists on $(0,T_{0}]$ for some computable $T_{0}>0$. Since $T_{\mathrm{crit}}$ for the large-$T$ bound is at most $19.2$ (for $d=14$), the two regimes overlap, covering all $T>0$.
Case 2: $d\in\{2,3,5,6,7,10,11,13\}$ (squarefree values below $14$). For these eight values we verify the inequality $\langle S_{L}^{2}\rangle_{T}>(S_{\zeta}^{*})^{2}/d$ for all $T>0$ by rigorous interval arithmetic using ARB at 512-bit precision, via three complementary bounds covering the entire half-line.

### 03 Mar 2026 v2

# Unconditional Density Bounds for Quadratic Norm-Form Energies via Lorentzian Spectral Weights Persistent Heuristics: Mini Monograph I

@@ Abstract.
For a real quadratic field $\mathbb{Q}(\sqrt{d})$, we study the norm-form energy $N=S_{\zeta}^{2}-d\cdot S_{L}^{2}$, where $S_{\zeta}$ and $S_{L}$ are Lorentzian-weighted zero sums with $w(\rho)=2/(\tfrac{1}{4}+\gamma^{2})$. We prove three main results. (1) Spacelike spectral data: $N<0$ unconditionally for all squarefree $d>1$, as a consequence of a low-lying zero dominance theorem proved via explicit zero-counting. (2) Effective density bound: at each verified truncation level $M$, $\operatorname{dens}\{N>0\}\leq 2\|f_{S_{L}^{(M)}}\|_{\infty}\cdot(W_{1}(\zeta)/\sqrt{d}+\epsilon_{M})$, established unconditionally via Jacobi–Anger resonance analysis. At fixed $M$ the bound is nontrivial only for sufficiently large $d$; the $O(1/\sqrt{d})$ rate requires $M$ to grow with $d$, which in turn requires a uniform density bound that we establish under a computationally verified finite-rank condition on the resonance lattice. (3) Exact asymptotic: under the computationally verified hypothesis that the infinite resonance lattice $\Lambda_{\infty}$ has finite rank (verified for $M\leq 20$, where rank =0=0), the sharp asymptotic $\operatorname{dens}\{N>0\}=C(d)/\sqrt{d}+o(1/\sqrt{d})$ holds. For $d=5$, $C(5)=2\,f_{S_{L}}(0)\cdot\mathbb{E}[|S_{\zeta}|]=0.1191$; the constant depends on $d$ through the zeros of $L(s,\chi_{d})$, and $C(d)=O(1/\log d)$ as $d\to\infty$.
Appendix F tabulates the first 1000 zeros at 70 decimal places for $L(s,\chi_{2})$, $L(s,\chi_{3})$, $L(s,\chi_{5})$, $L(s,\chi_{6})$, $L(s,\chi_{7})$, $L(s,\chi_{10})$, $L(s,\chi_{11})$, and $L(s,\chi_{13})$, all rigorously certified by ARB interval arithmetic.
@@ 1.2. Logical structure and the GSH barrier
Thm 4.17: Density vanishing Hypothesis 4.4(ii) $O(1/\sqrt{d})$ Jessen–Wintner
Thm 4.19: Euler product density nothing $O(1/\sqrt{d})$ $\mathbb{Q}$-indep. of $\log p$
@@ 1.4. Organization
Remark 1.1 (Purpose of Appendix F).
The proofs in this paper require only the zero data in Appendices BE: the first $200$ zeros of $L(s,\chi_{5})$ at $70$-digit precision, the first $20$ zeros of $L(s,\chi_{d})$ at $20$-digit precision for $d\in\{2,3,6,7,10,11,13\}$, and the first $60$ zeros of $\zeta(s)$. In the course of this work, however, we computed and certified the first 1,0001{,}000 zeros of $L(s,\chi_{d})$ at $70$ decimal places for all eight discriminants (8000 zeros in total); these extended tables are collected in the companion document Appendix F and are freely available as ancillary data files. To the best of our knowledge, no publicly available repository provides bulk high-precision zero ordinates for individual Dirichlet $L$-functions: the LMFDB [9] stores approximately $25$ zeros per character at $30$-digit precision (computed by Platt [13]), and the dataset of O’Bryant and Rechnitzer [14] provides zeros for all primitive characters of conductor up to $934$ at $12$-digit precision. We hope the present tables serve as a useful reference for researchers requiring certified zero data for quadratic $L$-functions.
@@ 2.1. The Lorentzian weight
Since every nontrivial zero of $\zeta(s)$ has $|\gamma|\geq 14.13>1/(2\sqrt{3})$, all zeta zeros are unconditionally $\gamma_{\mathrm{knw}}$ (critical-line membership verified). For $L$-function zeros, this paper uses two regimes. When $\beta=\tfrac{1}{2}$ is individually verified (by sign changes of the Hardy $Z$-function or ARB certification), the zero ordinate is denoted $\gamma^{\prime}_{\mathrm{knw}}$ and the on-line weight $w(\rho)=2/(\tfrac{1}{4}+\gamma^{\prime 2})$ is exact. When a zero’s existence is guaranteed by zero-counting (Theorem 3.2) but $\beta$ is not individually verified, the ordinate is denoted $\gamma^{\prime}_{\mathrm{unk}}$ and we use only the unconditional lower bound
$w(\rho)\;\geq\;w^{-}(\gamma^{\prime})\;:=\;\frac{2}{1+\gamma^{\prime 2}},$ (6)
valid for all $\beta\in(0,1)$. (Proof: writing $a=\beta(1-\beta)\in(0,\tfrac{1}{4}]$, one has $w(\rho)-2/(1+\gamma^{\prime 2})=2a(1-a+3\gamma^{\prime 2})/\bigl((a+\gamma^{\prime 2})^{2}+(1-2\beta)^{2}\gamma^{\prime 2}\bigr)(1+\gamma^{\prime 2})>0$.) Using the on-line weight $2/(\tfrac{1}{4}+\gamma^{\prime 2})$ for $\gamma^{\prime}_{\mathrm{unk}}$ zeros would implicitly assume the Riemann Hypothesis for $L(s,\chi_{d})$.
@@ 3.1. Low-lying zero dominance
We establish the inequality for all fundamental discriminants by combining Theorem 3.2 with verified computations for two small-conductor exceptions. Throughout, zeros are counted in the upper half-plane only, with the $\gamma^{\prime}_{\mathrm{knw}}$$\gamma^{\prime}_{\mathrm{unk}}$ convention and the lower bound $w^{-}$ (6) as defined above.
@@
where $\ell=\log(q\cdot 13/(2\pi))$. For $q=12$, this evaluates to $N(11,\chi_{d})\geq 3.80$; since $N(11,\chi_{d})$ is an integer, this implies at least four zeros with $\gamma^{\prime}_{\mathrm{unk}}\leq 11$, and the bound increases with $q$. Each contributes weight at least $w^{-}(11)=2/122=1/61$. Therefore
$S_{L}\geq 4\cdot w^{-}(11)=4/61=0.06557>0.04871=S_{\zeta}^{*}.$
Step 2: $d=5$ (conductor $q=5$). The BMOR bound gives only $N(11,\chi_{5})\geq 1.40$, which does not guarantee three zeros. We therefore verify the inequality directly. Computing the Hardy $Z$-function $Z_{\chi_{5}}(t)=e^{i\theta(t)}L(\tfrac{1}{2}+it,\chi_{5})$ and scanning for sign changes on $(0,12)$ yields two zeros below height $11$: $\gamma^{\prime}_{1,\mathrm{knw}}=6.6485$ and $\gamma^{\prime}_{2,\mathrm{knw}}=9.8314$, both verified to satisfy $|L(\tfrac{1}{2}+i\gamma^{\prime},\chi_{5})|<10^{-15}$. (Any additional zeros below height $11$, if they exist, would only increase $S_{L}$.) Therefore
$S_{L}\geq w(\gamma^{\prime}_{1,\mathrm{knw}})+w(\gamma^{\prime}_{2,\mathrm{knw}})=\frac{2}{0.25+44.20}+\frac{2}{0.25+96.66}=0.04499+0.02064=0.06563>S_{\zeta}^{*}.$
@@
$\gamma^{\prime}_{1,\mathrm{knw}}=4.8999739970070365010,\quad\gamma^{\prime}_{2,\mathrm{knw}}=7.6284288417693978341,\quad\gamma^{\prime}_{3,\mathrm{knw}}=10.806588163861712014,$
@@
$S_{L}\geq w(\gamma^{\prime}_{1,\mathrm{knw}})+w(\gamma^{\prime}_{2,\mathrm{knw}})+w(\gamma^{\prime}_{3,\mathrm{knw}})=\frac{2}{0.25+24.010}+\frac{2}{0.25+58.193}+\frac{2}{0.25+116.782}=0.08246+0.03421+0.01710=0.13377>S_{\zeta}^{*}.$
@@
The margin $(S_{L}-S_{\zeta}^{*})/S_{\zeta}^{*}$ is smallest in the generic case: for $q=12$, the BMOR bound guarantees at least four $\gamma^{\prime}_{\mathrm{unk}}$ zeros below height $11$, giving $S_{L}\geq 4\cdot w^{-}(11)=4/61=0.06557$, a margin of $34.6\%$ over $S_{\zeta}^{*}$. The margin is largest for $d=2$, where three $\gamma^{\prime}_{\mathrm{knw}}$ zeros give $S_{L}\geq 0.13375$, a margin of $174.6\%$.
@@
For most discriminants, a single directly computed zero suffices: $w(\gamma^{\prime}_{1,\mathrm{knw}})>S_{\zeta}^{*}$. The exception is $d=5$ ($\gamma^{\prime}_{1,\mathrm{knw}}=6.6485$), where $w(\gamma^{\prime}_{1,\mathrm{knw}})=0.04499<S_{\zeta}^{*}=0.04871$ and the second zero is required. In the generic $\gamma^{\prime}_{\mathrm{unk}}$ case (Step 1), at least four zeros are needed.
@@ 3.3. The negativity theorem for general characters
The proof follows the same structure as Theorem 3.1. For $q\geq 4$, Theorem 3.2 guarantees a $\gamma^{\prime}_{\mathrm{unk}}$ zero below height $T=10.5$, giving $q\cdot w^{-}(\gamma^{\prime}_{1,\mathrm{unk}})\geq 4\cdot 2/111.25=32/445=0.0719>S_{\zeta}^{*}$. For $q=3$, the BMOR bound at $T=15.7$ guarantees at least three $\gamma^{\prime}_{\mathrm{unk}}$ zeros below height $15.7$, so $q\cdot S_{L(\chi)}\geq q\cdot 3\cdot w^{-}(15.7)=3\cdot 3\cdot 2/(1+15.7^{2})=18/247.49=0.0727>S_{\zeta}^{*}$. ∎
@@
By (7), $S_{\zeta}\leq S_{\zeta}^{*}$. For $q\geq 4$, the BMOR bound guarantees a $\gamma^{\prime}_{\mathrm{unk}}$ zero below height $T(q)\leq 10.5<11$, so $S_{L(\chi)}\geq w^{-}(11)=1/61$. For $q=3$, the BMOR bound at $T=15.7$ gives $S_{L(\chi)}\geq 3\cdot w^{-}(15.7)=6/247.49>1/61$. In both cases, $q\cdot S_{L(\chi)}\geq 2q/122$, and $E_{\chi}(0)=S_{\zeta}-q\cdot S_{L(\chi)}\leq S_{\zeta}^{*}-2q/122$. Since $2\cdot 3/122=3/61=0.04918>S_{\zeta}^{*}=0.04871$, this is negative for all $q\geq 3$. ∎
@@ 3.5. Robustness
The negativity and spacelike results are robust to RH-failure for $\zeta$, and to RH-failure for $L(s,\chi_{d})$ in the presence of a Siegel zero. If the Riemann Hypothesis fails for $\zeta(s)$, off-line zeros have smaller Lorentzian weight by Proposition 2.2, decreasing $S_{\zeta}$. If RH fails for $L(s,\chi_{d})$ via a Siegel zero $\rho_{0}=\beta_{0}$ with $\beta_{0}$ near $1$, the weight $w(\rho_{0})=2/(\beta_{0}(1-\beta_{0}))$ diverges, increasing $S_{L}$. Both scenarios strengthen the inequalities.
@@ 4.1. Shifted zero sums
Each is a uniformly absolutely convergent cosine series ($\sum w(\rho)<\infty$), hence a uniform limit of trigonometric polynomials, and therefore uniformly almost periodic in the sense of Theorem 4.2. (If any zero has multiplicity $m>1$, group equal ordinates into a single frequency with coefficient equal to $m$ times the weight.) ∎
@@ 4.2. The Grand Simplicity Hypothesis
The unconditional results of this paper (Theorems 3.1, 3.7, 4.10, and 8.5) require neither sub-case.
@@ 4.3. Parseval identity
Assume Hypothesis 4.4 (i) (cross-function disjointness of zero ordinate sets) and simplicity of zeros within each function. Then the Cesàro mean of the norm-form energy satisfies
@@
so the disjointness-free identity gives a weaker (less negative) bound on $\langle N\rangle$. Without simplicity, the Parseval norms satisfy $\langle S_{\zeta}^{2}\rangle\geq\tfrac{1}{2}W_{2}(\zeta)$ (with equality if and only if all zeros are simple), so the exact formula requires simplicity while the negativity does not.
@@
By Theorem 4.2, applied to the grouped expansion (combining any repeated ordinates into a single frequency with coefficient equal to the total weight at that ordinate), $\langle S_{\zeta}(\delta)^{2}\rangle=\tfrac{1}{2}\sum_{\text{distinct }\gamma}c_{\gamma}^{2}$ and $\langle S_{L}(\delta)^{2}\rangle=\tfrac{1}{2}\sum_{\text{distinct }\gamma^{\prime}}{c^{\prime}_{\gamma^{\prime}}}^{2}$, where $c_{\gamma}=\sum_{\rho:\,\operatorname{Im}(\rho)=\gamma}w(\rho)$. Under simplicity of zeros (each ordinate appearing once), $c_{\gamma}=w(\gamma)$ and the sums reduce to $\tfrac{1}{2}W_{2}(\zeta)$ and $\tfrac{1}{2}W_{2}(L)$ respectively. The cross-term $\langle S_{\zeta}\cdot S_{L}\rangle=\tfrac{1}{2}\sum_{\gamma_{k}=\gamma_{j}^{\prime}}c_{\gamma_{k}}\,c^{\prime}_{\gamma_{j}^{\prime}}$ vanishes when the zero ordinate sets are disjoint. ∎
@@ 4.4. Cesàro spacelike theorem
For $k\geq 2$, set $a_{1}=w(\gamma_{1})=2/(\tfrac{1}{4}+\gamma_{1}^{2})$ (the largest zeta weight; an upper bound regardless of RH by critical line maximality, Proposition 2.2).
Remark 4.9 (Why the unconditional weight suffices).
The denominator penalty for not knowing $\beta=\tfrac{1}{2}$ is $3/4$: the unconditional weight uses $1+\gamma^{\prime 2}$ instead of $\tfrac{1}{4}+\gamma^{\prime 2}$. The crossover height where $w^{-}(T)=a_{1}$ is $T^{*}=\sqrt{\gamma_{1}^{2}-3/4}=14.108\ldots$, so any zero below height $14.108$ has $w^{-}>a_{1}$ unconditionally. Since $14<14.108$, the BMOR bound at $T=14$ lands safely below the crossover. This ensures $w^{-}(14)/a_{1}>1$, which makes $k=2$ the binding case. The actual $k=2$ verification then requires the count of BMOR-guaranteed zeros: a single zero does not suffice, since $w^{-}(14)=0.01015<S_{\zeta}^{*}=0.04871$, but $d\cdot n(d)$ zeros collectively overcome the deficit.
@@
$W_{k}(\zeta)=\sum_{j}a_{j}^{k}\leq a_{1}^{k-1}\sum_{j}a_{j}=a_{1}^{k-1}\cdot S_{\zeta}^{*}.$
Lower bound on $W_{k}(L)$: Theorem 3.2 at $T=14$ guarantees $n(d)$ zeros below height $14$ for every fundamental discriminant, each with unconditional weight at least $w^{-}(14)=2/197$. Therefore $W_{k}(L)\geq n(d)\cdot w^{-}(14)^{k}$.
Combining: the inequality $d\cdot W_{k}(L)>W_{k}(\zeta)$ follows from $d\cdot n(d)\cdot w^{-}(14)^{k}>a_{1}^{k-1}\cdot S_{\zeta}^{*}$. Since $w^{-}(14)/a_{1}=200.04/197=1.0154>1$ (see Remark 4.9), the left side grows faster in $k$ than the right; the binding constraint is $k=2$, which reduces to
$d\cdot n(d)>\frac{a_{1}\cdot S_{\zeta}^{*}}{w^{-}(14)^{2}}=4.725.$
Three cases exhaust all squarefree $d>1$. For $d=5$ ($q=5$): BMOR gives $n\geq 3$, so $d\cdot n\geq 15$. For $d=2$ ($q=8$): BMOR gives $n\geq 5$, so $d\cdot n\geq 10$. For all remaining $d$ ($q\geq 12$): BMOR gives $n\geq 7$ and $d\geq 3$, so $d\cdot n\geq 21$. The worst case is $d=2$, clearing the threshold by a factor of $2.12$. ∎
@@
for all $T>0$, where $c_{0}=(\pi-1)/(2\pi)$. For $q\geq 17$ and $T=11$, Theorem 3.2 guarantees at least five $\gamma^{\prime}_{\mathrm{unk}}$ zeros below height $11$, giving $W_{2}(L)\geq 5\,w^{-}(11)^{2}=5(8/488)^{2}=320/238144$. Therefore $c_{0}W_{2}(L)\geq 320(\pi-1)/(476288\pi)>4.58\times 10^{-4}$, while $(S_{\zeta}^{*})^{2}/d\leq(0.04871)^{2}/14<1.70\times 10^{-4}$, a margin of $2.7\times$ or better for all $d\geq 14$. For the worst-case five-$\gamma^{\prime}_{\mathrm{unk}}$-zero configuration (equally spaced in $[6,11]$), the interference constant evaluates to $\mathcal{I}=5.47\times 10^{-3}$, so the full Cesàro variance exceeds the threshold for all $T>3.1$.
@@
For small $T$: as $T\to 0^{+}$, $\langle S_{L}^{2}\rangle_{T}\to W_{1}(L)^{2}\geq(4\,w^{-}(11))^{2}=(4/61)^{2}=16/3721\approx 4.30\times 10^{-3}$, which exceeds $(S_{\zeta}^{*})^{2}/d\leq 1.70\times 10^{-4}$ by a factor of $25.4$. By the same Lipschitz bound used in Case 2 (with $|h^{\prime}(T)|\leq L$), the inequality persists on $(0,T_{0}]$ for some computable $T_{0}>0$. Since $T_{\mathrm{crit}}$ for the large-$T$ bound is at most $19.0$ (for $d=14$), the two regimes overlap, covering all $T>0$.
Case 2: $d\in\{2,3,5,6,7,10,11,13\}$ (squarefree values below $14$). For these eight values we verify the inequality $\langle S_{L}^{2}\rangle_{T}>(S_{\zeta}^{*})^{2}/d$ for all $T>0$ by rigorous interval arithmetic using ARB at 512-bit precision, with all zero ordinates $\gamma^{\prime}_{\mathrm{knw}}$ individually verified on the critical line. Three complementary bounds cover the entire half-line.